OpenStudy (anonymous):

A beam of protons is moving toward a target in a particle accelerator. This beam constitutes a current whose value is 0.50 μA. How many protons strike the target in 15 seconds?

OpenStudy (anonymous):

a) 4.7×1013 * b) 1.6 × 10-19 c) 5.3 × 10-25 d) 4.7 × 1012 e) none of the above

OpenStudy (anonymous):

a

OpenStudy (anonymous):

i = 5*10 -7 A . q= 5 *15 *10 -7 coulomb = 75*10 -7 C in 15 sec is incident but 1 proton has q = 1.6 *10 -19 C hence for above charge nos of protons are 75 *10 -7) /1.6*10 -19. =4.68 * 10 13 = 4.7 10 13 (a)